# Cross-Modal Conditioned Reconstruction for Language-guided Medical Image Segmentation

### v1

@@ 3.3.1 Conditioned Vision Reconstruction
where $\sqrt{D}$ is the feature dimension and $\odot$ denotes the Hadamard product. Then we utilize the mean squared error to measure the similarity between the reconstructed $\hat{V}$ and the original visual features $V$:
@@
3.3.2 Conditioned Language Reconstruction
To reconstruct textual features $V$ from visual features, we perform a similar operation as described in CVRLABEL:CVR. Let’s define the reconstructed text features as $\hat{E}=\{\hat{e}\}_{i=1}^{L}$ and $E_{m}$ denotes the masked textual features. The process of conditioned language reconstruction can be expressed as:
@@ 3.4 Training and Inference
$\mathcal{L}\!=\!\lambda_{1}(\mathcal{L}_{\mathrm{T}2\mathrm{V}}+\mathcal{L}_{\mathrm{V}2\mathrm{T}})+\lambda_{2}\mathcal{L}_{\mathrm{CCL}}+\lambda_{3}(\mathcal{L}_{\mathrm{Dice}}+\mathcal{L}_{\mathrm{CE}}),$ (10)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are trade-off hyper-parameters.
Inference. As shown in Fig. 2, the orange arrows denote the inferring path. We only need to keep the visual encoder, visual decoder, and text-to-vision encoder in the CI module. In contrast to previous work (e.g., [1], [36]), our RecLMIS eliminates the need for complex multi-level interaction, which significantly accelerates inference, as detailed in Sec.4.4.7.
@@ 4.1.1 QaTa-COV19 Dataset
Method Backbone Text Parameters Flops QaTa-COV19 MosMedData+
Dice(%$\uparrow$ mIoU(%$\uparrow$ Dice(%$\uparrow$ mIoU(%$\uparrow$
U-Net [17] CNN 14.8 50.3 79.02 69.46 64.60 50.73
U-Net++ [18] CNN 74.5 94.6 79.62 70.25 71.75 58.39
AttUNet [47] CNN 34.9 101.9 79.31 70.04 66.34 52.82
nnUNet [20] CNN 19.1 412.7 80.42 70.81 72.59 60.36
TransUNet [21] Hybrid 105.0 56.7 78.63 69.13 71.24 58.44
Swin-Unet [22] Hybrid 82.3 67.3 78.07 68.34 63.29 50.19
UCTransNets [48] Hybrid 65.6 63.2 79.15 69.60 65.90 52.69
ConVIRT [49] CNN 35.2 44.6 79.72 70.58 72.06 59.73
TGANet [16] CNN 19.8 41.9 79.87 70.75 71.81 59.28
CLIP [7] Hybrid 87.0 105.3 79.81 70.66 71.97 59.64
GLoRIA [50] Hybrid 45.6 60.8 79.94 70.68 72.42 60.18
ViLT [51] Hybrid 87.4 55.9 79.63 70.12 72.36 60.15
LAVT [36] Hybrid 118.6 83.8 79.28 69.89 73.29 60.41
LViT [1] Hybrid 29.7 54.1 83.66 75.11 74.57 61.33
RecLMIS CNN 23.7 24.1 85.22 77.00 77.48 65.07
$\Delta$20.2% -55.5% +1.56 +1.89 +2.91 +3.74
@@ 4.2 Implementation Details
We implemented our method in PyTorch [52] and the training devices were four NVIDIA GeForce RTX 3090 with 24GB video memory capacity per card. The initial learning rate was set to 3e-4, and we utilized cosine annealing as the scheduler. The default batch size was 32 and the resolution of the image was $224\!\times\!224$. We trained the networks on GPUs with 200 epochs for QaTa-COV19 [46] dataset and 500 epochs for MosMedData+ [2, 3, 1]. The visual mask ratio $\alpha_{v}$ and textual mask ratio $\alpha_{t}$ were 0.5 and 0.3, respectively. The number of conditioned transformer layers was three. Besides, we applied Gaussian blur with random sigma, contrast distortion, saturation distortion, and hue distortion to augment the datasets. The $\lambda_{1},\lambda_{2},\lambda_{3}$ were 1, 0.2, and 2.5, respectively.
@@ 4.3 Comparison with State-Of-The-Art
QaTa-COV19 and MosMedData+. In Table 1, we evaluate RecLMIS against the other segmentation methods on the QaTa-COV19 [46] and MosMedData+ [2]. Compared to the previous optimal model LViT [1], our rec surpasses it by 1.33 mIoU and 1.18 Dice on QaTa-COV19 and 2.91 mIoU and 3.74 Dice on MosMedData+. More encouragingly, RecLMIS exhibits a relative reduction of 55.5% in computational cost and a relative decrease of 20.2% in parameter count. These results demonstrate the superiority and efficiency of RecLMIS. Moreover, RecLMIS improves the Dice score by 4.42 compared to the suboptimal nnUNet [20] without text on QaTa-COV19, which indicates that introducing medical notes can improve model performance effectively in chest X-ray radiograph segmentation.
@@ 4.4 Ablation Study
Ablation experiments on the challenging QaTa-COV19 [46] dataset are conducted to substantiate the effectiveness of RecLMIS, which is explored in the following seven aspects.
@@ 4.4.1 Effect of Proposed Components
We consider the U-Net [17] model with the PLAM [1] module but without textual prompts as our baseline. As shown in the last two rows of Table 2, mIoU is significantly improved when we introduce the Conditioned Interaction (CI) module, thanks to the textual information that provides effective guidance for segmentation. By comparing the first three rows, the inclusion of CLR and CVR results in an improvement of 0.3 and 0.8 absolute points in mIoU, respectively. This indicates that CVR and CLR play a significant role in facilitating effective alignment between the visual and textual modalities. Our full model achieves the best performance and outperforms the baseline model by 6.44 mIoU.
@@ 4.4.2 Effect of Condition
As described in Sec. 3.2 and Sec. 3.3, we emphasize the interdependence and influence between contrastive learning and reconstruction processes in both textual and visual features. Therefore, in this section, we explore the effects introduced by condition constraints. As shown in Table 3, the mIoU metric decreases when we remove the condition constraints $w_{v}$ and $w_{e}$ in any module. When we remove all conditioned constraints, the performance drops by 0.87 mIoU.
4.4.3 Ablation Study on Mask Strategy.
As illustrated in Table 4, we have explored two masking mechanisms: one involves random word selection for text and random patch selection for images, while the other relies on the obtained $W_{woi}$ and $W_{poi}$ as probability distributions to generate masks as Sec. 3.3 mentioned. Employing the probability distribution-based method enables the model to autonomously learn the critical information locations in both images and text, which is more conducive to cross-modal learning.
@@ 4.4.4 Effect of Masking Ratio
As Sec. 4.4.3 described, the probability distribution-based method works better, so we explore the impact of different text and image masking ratios built upon this foundation. Specifically, in Fig. 5, it can be observed that when the masking ratios for $\alpha_{t}$ and $\alpha_{v}$ are set to 0.3 and 0.5 respectively, the probability-based approach yields the best results. This reveals that images contain considerable redundant information, leading to better outcomes with larger masking ratios, consistent with the conclusions drawn in MAE. However, unlike MAE, our masking is not randomly selected but learned through the model. Hence, setting the image masking ratio to 0.9 in our experiments yields suboptimal results.
@@ 4.4.5 Effect of the mechanism of cross-modal interaction
Figure 7: Visualization results. The left-hand image represents the visualization results on the test set of QaTa-COV19, while the right-hand image represents the visualization results on the test set of MosMedData+. Green, red, and blue, indicate true positive, false negative, and false positive pixels, respectively. Orange font signifies words that carry instructive significance for the segmentation results. Best viewed in color, and further enhance the view by zooming in.
Figure 8: Visualization of the class activation heatmaps. The left-hand image represents the visualization results on the test set of QaTa-COV19, while the right-hand image represents the visualization results on the test set of MosMedData+. The column titled “medical note” denotes the input textual prompt, while the column labeled “Image” signifies the input image. The column labeled “GT” represents the ground truth segmentation target. Columns labeled “Baseline”, “LViT” and “Ours” depict the class activation heatmaps computed by these three models based on the gradient of the first layer of the decoder elucidating the efficacy of attention in the models. Best viewed in color, and further enhance the view by zooming in.
@@ 4.4.6 Effect of the Number of Reconstruction Layer
4.4.7 Efficiency of our method
Table 5: The efficiency of the different models on the QaTa-COV19 dataset.
Method Dice(%$\uparrow$ mIoU(%$\uparrow$
Inf. Time
(ms/img)
LAVT [36] 79.28 69.89 170.3
TGANet [16] 79.87 70.75 51.5
LViT [1] 83.66 75.11 37.8
RecLMIS 84.84 76.44 20.3
In Table 5, we compute the inference time using an NVIDIA GeForce RTX 3090 on the QaTa-COV19 [46] dataset. Since we have discarded complex multi-level cross-modal interaction methods (e.g.LViT [1]), as well as CVR and CLR can be removed during inference, our approach requires just 20.3ms to process a single image, nearly twice as fast as LViT [1]. This underscores the efficiency of our design.
4.5 Qualitative Analysis
As shown in Fig. 7, the RecLMIS we proposed can accurately segment, regardless of size or shape. Whether on the QaTa-COV19 [46], MosMedData+ [2, 3, 1], our model has outperformed LViT [1]. For interpretability, we employ Grad-CAM [53] to illustrate the class activation heatmaps computed by these three models based on the gradient of the first layer of the decoder. As shown in Fig. 8, we find that RecLMIS can achieve improved alignment between visual and textual features, compared to LViT [1], RecLMIS focuses on regions that align more closely with the semantics of the text.

### v2

@@ 3.3 Conditioned Vision Reconstruction
where $\sqrt{D}$ is the feature dimension and $\odot$ denotes the Hadamard product.
Then we utilize the mean squared error to measure the similarity between the reconstructed $\hat{V}$ and the original visual features $V$:
@@
3.4 Conditioned Language Reconstruction
Using masked text to reconstruct raw text has been well-developed in method BERT [6]. Many works [4, 9] have adopted such approaches to enhance image-to-text representation capabilities. These methods fall under the category of Fig. 2(a), which involves masking and reconstructing the text itself. While such schemes can improve the model’s vision-to-text representation capabilities, reconstructing the original text includes redundant words, introducing noise. We aim to introduce an auxiliary module that, without modifying the original model structure, utilizes visual features $V$ to reconstruct the masked keywords in the text features $E_{m}$, thereby achieving a better understanding from vision to text.
As shown in Fig. 5(b), to reconstruct masked textual features from visual features, we perform a similar masking operation as described in CVR 3.3. Let’s define the reconstructed text features as $\hat{E}=\{\hat{e}\}_{i=1}^{L}$ and $E_{m}$ denotes the masked textual features. The process of conditioned language reconstruction can be expressed as:
@@ 3.4 Conditioned Language Reconstruction
Table 2: The efficiency of the different models on the QaTa-COV19 dataset. ’Cov.’ and ’Inf.’ denotes ’Convergence’ and ’Inference’, respectively.
Method Training Phase Inference Phase
Params (M) FLOPs (G) Con. Time (hours) Params (M) FLOPs (G) Inf. Time (ms/img)
LAVT [7] 118.6 83.8 - 118.6 83.8 170.3
LViT [1] 29.7 54.1 16 29.7 54.1 37.8
DMMI [9] 114.6 63.3 8.6 114.6 63.3 67.4
RefSegformer [41] 195.0 103.6 13.5 195.0 103.6 75.6
RecLMIS(ours) 74.2 36.62 7.5 23.7 24.1 20.3
@@ 3.5 Training and Inference
$\mathcal{L}\!=\!\lambda_{1}\mathcal{L}_{\mathrm{V}2\mathrm{T}}+\lambda_{2}\mathcal{L}_{\mathrm{T}2\mathrm{V}}+\lambda_{3}\mathcal{L}_{\mathrm{CCL}}+\lambda_{4}(\mathcal{L}_{\mathrm{Dice}}+\mathcal{L}_{\mathrm{CE}}),$ (11)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are trade-off hyper-parameters.
Inference. As shown in Fig. 3, the orange arrows denote the inferring path. We only need to keep the visual encoder, visual decoder, and text-to-vision encoder in the CI module. In contrast to previous work (e.g., [1], [7]), our RecLMIS eliminates the need for complex multi-level interaction, which significantly accelerates inference, as detailed in Sec.4.3.2).
@@ 4.2 Implementation Details
We implemented our method in PyTorch [59] and the training devices were four NVIDIA GeForce RTX 3090 with 24GB video memory capacity per card. The initial learning rate was set to 3e-4, and we utilized cosine annealing as the scheduler. The default batch size was 32 and the resolution of the image was $224\!\times\!224$. We trained the networks on GPUs with 200 epochs for QaTa-COV19 [58] dataset and 500 epochs for MosMedData+ [2, 3, 1]. The visual mask ratio $\alpha_{v}$ and textual mask ratio $\alpha_{t}$ were 0.5 and 0.3, respectively. The number of conditioned transformer layers was three. Besides, we applied Gaussian blur with random sigma, contrast distortion, saturation distortion, and hue distortion to augment the datasets. The $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ were 1, 1, 0.2, and 5, respectively.
@@ 4.3 Comparison with State-Of-The-Art
4.3.1 Comparison on QaTa-COV19 and MosMedData+ Dataset
Figure 6: The visualization of the main ablation experiments Sec. 4.4.1). The sequence numbers correspond to those in Table 3. Green, red, and blue, indicate true positive, false negative, and false positive pixels, respectively.
In Table 1, we evaluate RecLMIS against the other segmentation methods on the QaTa-COV19 [58] and MosMedData+ [2]. Compared to the previous optimal model LViT [1], our rec surpasses it by 1.33 mIoU and 1.18 Dice on QaTa-COV19 and 2.91 mIoU and 3.74 Dice on MosMedData+. More encouragingly, RecLMIS exhibits a relative reduction of 55.5% in computational cost and a relative decrease of 20.2% in parameter count. These results demonstrate the superiority and efficiency of RecLMIS. Moreover, RecLMIS improves the Dice score by 4.42 compared to the suboptimal nnUNet [25] without text on QaTa-COV19 [58], which indicates that introducing medical notes can improve model performance effectively in chest X-ray radiograph segmentation.
As shown in Fig. 9, we conducted a visualization analysis of the main ablation experiments. The RecLMIS we proposed can accurately segment, regardless of size or shape. Whether on the QaTa-COV19 [58], MosMedData+ [2, 3, 1], our model has outperformed LViT [1].
4.3.2 Efficiency comparison
In Table 2, we compute the inference time using an NVIDIA GeForce RTX 3090 on the QaTa-COV19 [58] dataset. During the training phase, our RecLMIS demonstrates balanced performance. Despite a higher parameter count in RecLMIS training than other methods, both the FLOPs and training convergence time were significantly curtailed, resting at a mere 7.5 hours, less than half of the 16 hours required by LViT [1]. Upon entering the inference phase, RecLMIS continued to showcase superiority. Since we have discarded complex multi-level cross-modal interaction methods (e.g., LViT [1], LAVT [7]), as well as CVR and CLR can be removed during inference, our method achieves faster processing speeds during the inference phrase. Our model significantly reduced both the parameter count and FLOPs to 23.7M and 24.1G respectively, representing a reduction of 20.2% and 55.5% compared to LViT [1]. Moreover, the inference time of RecLMIS was clocked at 20.3 milliseconds per image, a speed increase of 46.3% compared to LViT [1], markedly outpacing both LAVT [7] and other methods [9, 41].
Figure 9: The visualization of the main comparison. The left-hand image represents the visualization results on the test set of QaTa-COV19, while the right-hand image represents the visualization results on the test set of MosMedData+. Green, red, and blue, indicate true positive, false negative, and false positive pixels, respectively. Orange font signifies words that carry instructive significance for the segmentation results. Best viewed in color, and further enhance the view by zooming in.
Figure 10: The visualization of the class activation heatmaps. The left-hand image represents the visualization results on the test set of QaTa-COV19, while the right-hand image represents the visualization results on the test set of MosMedData+. The column titled “medical note” denotes the input textual prompt, while the column labeled “Image” signifies the input image. The column labeled “GT” represents the ground truth segmentation target. Columns labeled “Baseline”, “LViT” and “Ours” depict the class activation heatmaps computed by these three models based on the gradient of the first layer of the decoder elucidating the efficacy of attention in the models. Best viewed in color, and further enhance the view by zooming in.
In summary, RecLMIS not only outperforms other methods in mIoU metrics but also significantly curtails computational resource consumption and inference time, thus underscoring its potential and advantages in practical applications.
@@ 4.4 Ablation Study
Ablation experiments on the challenging QaTa-COV19 [58] dataset are conducted to substantiate the effectiveness of RecLMIS, which is explored in the following seven aspects. For all ablation experiments except those figures (Fig 7, Fig 8, Fig 12), the text masking ratio and the visual masking ratio are both 0.3 in this section.
@@ 4.4.1 Effect of Proposed Components
We consider the U-Net [22] model with the PLAM [1] module but without textual prompts as our baseline. As shown in the last two rows of Table 3, mIoU is significantly improved when we introduce the CI module, thanks to the textual information that provides effective guidance for segmentation. By comparing the first three rows, the inclusion of CLR and CVR results in an improvement of 0.3 and 0.8 absolute points in mIoU, respectively. This indicates that CVR and CLR play a significant role in facilitating effective alignment between the visual and textual modalities. Our full model achieves the best performance and outperforms the baseline model by 6.44 mIoU.
As shown in Fig. 6, we conduct a visualization analysis of the main ablation experiments. It is evident that without the CI module incorporating textual information, there are many instances of incorrect positional predictions. When the CVR and CLR modules are introduced, the accurate boundary predictions are further refined.
@@ 4.4.2 Effect of Condition
As described in Sec. 3.2, Sec. 3.3, and Sec. 3.4, we emphasize the interdependence and influence between contrastive learning and reconstruction processes in both textual and visual features. Therefore, in this section, we explore the effects introduced by condition constraints. As shown in Table 4, the mIoU metric decreases when we remove the condition constraints $w_{v}$ and $w_{e}$ in any module. When we remove all conditioned constraints, the performance drops by 0.87 mIoU.
4.4.3 Ablation Study on Mask Strategy
We explore two masking mechanisms: one involves random text token and image patch selection, while the other relies on the obtained $W_{woi}$ and $W_{poi}$ as probability distributions to generate masks as Sec. 3.3 mentioned. Employing the probability distribution-based method enables the model to autonomously learn the critical information locations in both images and text, which is more conducive to cross-modal learning.
@@ 4.4.4 Effect of Masking Ratio
As Sec. 3.3 described, the probability distribution-based method works better, so we explore the impact of different text and image masking ratios built upon this foundation. Specifically, in Fig. 8, it can be observed that when the masking ratios for $\alpha_{t}$ and $\alpha_{v}$ are set to 0.3 and 0.5 respectively, the probability-based approach yields the best results. This reveals that images contain considerable redundant information, leading to better outcomes with larger masking ratios, consistent with the conclusions drawn in MAE. However, unlike MAE, our masking is not randomly selected but learned through the model. Hence, setting the image masking ratio to 0.9 in our experiments yields suboptimal results.
@@ 4.4.6 Effect of the Number of Reconstruction Layer
4.4.7 Effect of loss function weights and learning rate
Figure 12: Effect of $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ and the learning rate. Different background colors represent the impact brought by the variation of specific hyperparameters. the text masking ratio is 0.3 and the visual masking ratio is 0.5.
We conducted ablation studies targeting $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ and the learning rate in Fig. 12. We observed the most significant performance decline when $\lambda_{1}$ and $\lambda_{3}$ were increased to 1.5 and 0.5, respectively. We believe this is due to the large auxiliary losses severely disrupting the convergence of the primary segmentation loss. Overall, despite these hyperparameter fluctuations, the minimum performance remained above LViT’s 75.11.
Figure 13: The visualization of key variables ($W_{poi},W_{woi},V$, and $V^{\prime}$).
Figure 14: The visualization of the impact of text prompts on the results. Green, red, and blue, indicate true positive, false negative, and false positive pixels, respectively.
5 Qualitative Analysis
Quantitative analysis of the model comparison. For interpretability, we employ Grad-CAM [60] to illustrate the class activation heatmaps computed by these three models based on the gradient of the first layer of the decoder. As shown in Fig. 10, we find that RecLMIS can achieve improved alignment between visual and textual features, compared to LViT [1], RecLMIS focuses on regions that align more closely with the semantics of the text.
Quantitative analysis of the CI module. As depicted in the final two columns of Fig. 13, the image features extracted by the visual encoder are highly abstract, devoid of semantically matching textual features. However, after integrating the CI module, the features have been significantly enhanced to align with the semantic attributes of the given text prompt. This signifies that the model is paying more attention to the areas described by the text. Additionally, for the fused feature $V^{\prime}$ in Fig. 13(e), we observed that its focus areas are consistent with the regions indicated by the high-weight words in $W_{woi}$. This demonstrates that our visual and textual features are well-aligned.
Quantitative analysis of $W_{poi}$. For $W_{poi}$ in Fig. 13(b), the model adaptively focuses on the boundary regions between the foreground and background in the images. Notably, the lung boundary regions have higher weights. We believe this is because, in segmentation tasks, the pixels at the foreground-background boundaries are hard pixel cases, meaning they are the most challenging for the model to identify correctly. Our adaptive mask mechanism captures these hard pixel cases, masking and reconstructing them to better recognize these regions, demonstrating the effectiveness of our CVR module.
Quantitative analysis of $W_{woi}$. For $W_{woi}$ in Fig. 13(c), the model pays more attention to critical positional words like ’upper’, ’left’, ’middle’, and ’lower’. Non-positional keywords, such as ’pulmonary’, ’area’, and ’lung’, have relatively lower weights. This also proves the efficacy of our CLR module.
Quantitative analysis of text-to-image understanding. In Fig. 14, we remove the text prompt indicating left lung infection and used a new prompt. The results in the last column show that the model doesn’t predict the infection area in the left lung, but only in the right lung. This demonstrates that the model is well-aligned with the text, indicating it can understand the meaning of the text in the context of the image.
@@ 6 Conclusion
7 Limitations
Although our work represents the most advanced research in the Language-guided Medical Image Segmentation (LMIS) field, there are still some limitations. We identify two main shortcomings in this work: 1. We find that the performance on the MosMedData+ test set is lower than on the QaTa-COV19 test set primarily because the former contains more small target samples. The model performs poorly on small targets, indicating room for improvement in this area. 2. Although word order has minimal impact on the model, we observed that synonyms do affect the output. We believe that employing large language models (LLMs) to increase data diversity and developing an image-sentence level contrastive loss can alleviate these issues in future work.